# Local Spherical Collapsing Box in Athena++: Numerical Implementation and Benchmark Tests

### v1

@@
We implement a local model for a spherical collapsing/expanding gas cloud into the Athena++ magnetohydrodynamic code. This local model consists of a Cartesian periodic box with time-dependent geometry. We present a series of benchmark test problems, including non-linear solutions and linear perturbations of the local model, confirming the code’s desired performance. During a spherical collapse, a horizontal shear flow is amplified, corresponding to angular momentum conservation of zonal flows in the global problem; wave speed and amplitude of sound waves increase in the local frame, due to the reduction in the characteristic length scale of the box, which can lead to an anisotropic effective sound speed in the local box. Our code conserves both mass and momentum to machine precision. This numerical implementation of the local model has potential applications to the study of local physics and hydrodynamic instabilities during protostellar collapse, providing a powerful framework for better understanding the earliest stages of star and planet formation.
@@ 1 Introduction
Star formation has been studied over decades both theoretically and numerically on large spatial scales, key motivation being to determine how global redistribution of angular momentum results in the formation of stars, disks, or jets and set the initial conditions for planet formation (Larson 1969; Penston 1969 and e.g. Pattle et al. 2023; Pineda et al. 2023; Tsukamoto et al. 2023 for recent reviews). The field is stimulated by instruments capable of resolving the largest scales of these young stellar systems in ever finer detail in several wavelength domains, such as the Atacama Large Millimeter Array (ALMA), the Very Large Telescope (VLT) or the James Webb Space Telescope (JWST). The importance of these questions has perhaps overshadowed that of the physical processes taking place simultaneously on a small scale. Is it possible, for example, to start concentrating dust as soon as the cloud collapses, thereby promoting highly efficient planet formation later on?
@@
The local box formalism provides a simple treatment of the hydrodynamics of a spherically symmetric system undergoing contraction. The global effects of curvature and acceleration due to collapse are consistently incorporated into the local treatment of the problem by source terms that can be time-dependent, while allowing for the adoption of periodic boundaries as the system is spatially homogeneous. Numerical simulations using the local approximation approach play an important role in the study of local gas dynamics and instabilities in astrophysics. The best-known example is the rediscovery of the magnetorotational instability (MRI) with the local shear box approximation, which was first applied to magnetohydrodynamic (MHD) simulations by Hawley et al. (1995). Local models have been used for (isotropic) cosmological collapses (Robertson & Goldreich, 2012). Anisotropy in collapse/expansion was taken into account for the study of starless molecular core collapse (Toci et al., 2018), solar winds (Velli et al., 1992; Grappin et al., 1993; Grappin & Velli, 1996; Tenerani & Velli, 2017; Shi et al., 2020; Squire et al., 2020) (among which Tenerani & Velli (2017) also consider non-uniform expansion), for anisotropy-driven plasma kinetic instabilities (Sironi & Narayan, 2015; Bott et al., 2021), and for the cosmic ray pressure anisotropy instability (Sun & Bai, 2023).
@@
Recently, Lynch & Laibe (2023) (hereafter LL23) developed a new local model for the spherical collapse/expansion problem, allowing for a non-uniform collapse/expansion. This models considers a radially varying flow as in Tenerani & Velli (2017), but also allows for the time dependence of the background flow. It also uses a different treatment of pressure gradient than the expanding box models, in an approach that is more similar to the distorted shearing box models of Ogilvie & Latter (2013); Ogilvie & Barker (2014). This avoids spurious instabilities from the inclusion of formally subdominant terms (Latter & Papaloizou, 2017), allows for the application of periodic or shear-periodic boundary conditions, with a Cartesian like geometry. In this study, we incorporate LL23’s model into Athena++, a high order Godunov MHD code (Stone et al., 2020). Other local models have already been implemented in Athena++, including Squire et al. (2020) for (uniformly) expanding box, and MHD-PIC (Sun & Bai, 2023), also allowing for an anisotropic collapse/expansion. However, most of these previous numerical implementations focus on the study of MHD properties and magnetized waves, in the context of stellar winds or cosmic rays, and the hydrodynamic properties of the expanding/collapsing box have not been comprehensively tested so far. Here, we aim to incorporate the local model of LL23, with benchmark tests to thoroughly investigate and validate the hydrodynamic properties therein.
@@
This article is organized as follows. We first describe the formulation of the local model for collapsing box in §2. In §3, We describe the numerical implementation, especially the modification of the Riemann solver in order to incorporate the local model. We then conduct series of test cases including nonlinear solutions and linear perturbations in §4, as well as numerical studies. We further discuss possible applications and perspectives, and give future perspectives in §5. We summarize and conclude in §6.
@@
Figure 1: Cartoon illustration of the local collapsing box model. On the left, the global geometry depicts a domain moving radially due to background flow. The volume of the domain varies as it approaches the center of the spherical cloud, leading to changes in density. A compressing domain leads to a density increase, as shown by the progressively darker blue color in the sequence. The horizontal domain size is proportional to $\mathcal{R}\left(t\right)$, establishing the horizontal length scale in the local model (middle panel). The vertical length scale, $L_{z}\left(t\right)$, is determined by the distance between the domain’s inner and outer radial boundaries, which is set by the background flow. Inside each box, a waveform signal illustrates how certain physical quantities evolve in the local model. The width of the waveform represents quantities that remains fixed in the local frame (e.g. wavelength). The amplitude of the waveform represents specific physical quantities that are expected to be constant in the global geometry throughout the collapse (e.g. angular momentum, isothermal physical sound speed). In the local reference frame (right), the corresponding quantities such as the shear flow velocity or the effective sound speed are amplified as the collapse proceeds and the length scale shrinks in the domain, while quantities such as the wavelength remain the same. This phenomenon is visually demonstrated by the vertically stretched waveforms in the sequence. For a more accurate and detailed understanding of this process, one should refer to the equations in §2 .
@@ 2 Equations of the Local Model
Consider an axisymmetric spherical gas cloud collapsing or expanding with a purely radial velocity $U_{bg}(r,t)$ that varies with both radius $r$ and time $t$. Let’s now consider a local neighborhood of a reference point located at radius $\mathcal{R}(t)$ in the spherical gas cloud, described by a local Cartesian coordinate system comoving with the background flow. The local coordinates are given by
@@
where $\phi$ and $\theta$ are the longitude and latitude on the reference sphere of radius $\mathcal{R}$, $r-\mathcal{R}(t)$ is the local radial distance relative to the global $\mathcal{R}(t)$ of the reference frame. Note that $x$, $y$, $z$ have dimension of angles in this formalism, and that they denote two azimuthal and radial directions respectively. $L_{z}$ is a characteristic vertical length scale which is set by the background flow profile, which can be approximated by
@@
The flow within the local box model represents local, nonlinear perturbations to the background flow. Following the derivation in §2 of LL23, in the local frame, the continuity, momentum, and energy equations are given in the stretched Cartesian coordinate system by
@@
where $\rho$ is the density of the gas, $v_{i}$ is the relative velocity field in the local frame  [^1] The velocities here are the contravariant components of the velocity. In this article, we use subscript indices to denote the spatial dimensions (e.g., $v_{i}$) for convenience, as opposed to superscript indices (e.g., $v^{i}$) as is conventionally used. We also apply Einstein summation throughout this paper, where repeated indices (subscript or superscript) are implicitly summed over., $p$ is the pressure of the gas, and $\mathcal{E}_{rell}$ is the energy density of the relative motion in this system, (re-)defined as
$\mathcal{E}_{\text{rell }}=\frac{1}{2}L_{z}^{2}v_{z}^{2}+\frac{1}{2}\mathcal{R}^{2}(t)\left(v_{x}^{2}+v_{y}^{2}\right)+\varepsilon$ (10)
which is the sum of the kinetic energy density of relative motion in the local frame, $\mathcal{E}_{k,\text{rell }}\equiv\frac{1}{2}L_{z}^{2}v_{z}^{2}+\frac{1}{2}\mathcal{R}^{2}(t)\left(v_{x}^{2}+v_{y}^{2}\right)$, and the specific internal energy $\varepsilon$ determined by the equation of state. For an isothermal equation of state, $p=\rho c_{s}^{2}$, where $c_{s}$ is the sound speed, and the energy equation (9) is dropped from the system. For an adiabatic equation of state, $\varepsilon=\frac{p}{\rho(\gamma-1)}$ , where $\gamma$ is the ratio of specific heats, assumed to be constant.
The right-hand side of Eqs. 68 can be interpreted as an anisotropic sound speed in the local coordinate. We define two rescaled sound speeds on horizontal and vertical directions separately: $c_{\mathrm{s},\mathrm{H}}\equiv c_{\mathrm{s}}/\mathcal{R}$, and $c_{\mathrm{s},\mathrm{V}}\equiv c_{\mathrm{s}}/L_{z}$.
Eqs. 59 can be written in a compact form
@@
where the vector $\mathbf{q}$ of conservative variables is defined according to
@@
$\mathbf{f},\mathbf{g}$, and $\mathbf{h}$ are the fluxes of the associated conservative variables in the $x$, $y$, and $z$-directions respectively
@@
The vector of source terms, denoted $\mathbf{S}$, is
@@
Again, for an isothermal equation of state, the fifth component of the vectors $\mathbf{q}$, $\mathbf{f}$, $\mathbf{g}$, $\mathbf{h}$, and $\mathbf{S}$ are dropped, as the energy equation is not solved for this case.
A more intuitive illustration of the local model geometry is presented in Figure 1. As the collapse proceeds, the local domain is compressed in volume, leading to increase in gas density. This is reflected in the source term in the continuity equation (5). The angular momentum conservation law leads to amplification of horizontal flow velocity in the local frame during the collapse, which is represented by the source terms in the momentum equations on the horizontal directions (Eqs. 6 and 7). These features along with other characteristic behaviors of the local model are further illustrated and tested in §4.
@@ 3 Numerical Implementation
We now translate the local model of § 2 into a Godunov scheme, and implement it in the numerical code Athena++. Finite volume discretization of the compact form of the hydrodynamic equations Eq.(11) gives the time integration step as a set of ordinary differential equations
@@
where subscript indices $(i,j,k)$ denotes numerical cells centered at position $(x_{i},y_{j},z_{k})$, superscripts $n$ denotes the index for discrete time, $\delta x$, $\delta y$, $\delta z$ are discrete grid sizes, $\delta t$ is the time step. The vectors $\mathbf{q}_{i,j,k}^{n}$ and $\mathbf{S}_{i,j,k}^{n}$ correspond to volume-averaged $\mathbf{q}$ and $\mathbf{S}$ over the numerical cells. The vectors $\mathbf{f}_{i+1/2,j,k}^{n+1/2}$, $\mathbf{g}_{i+1/2,j,k}^{n+1/2}$ and $\mathbf{h}_{i+1/2,j,k}^{n+1/2}$ correspond to time-averaged $\mathbf{f}$, $\mathbf{g}$, and $\mathbf{h}$ respectively. The half-integer subscripts on the flux vectors refer to the edges of computational cells, while the half-integer superscripts denote time average between $t^{n}$ and $t^{n+1}$. $\mathbf{\tilde{q}}_{i,j,k}^{n}$ denotes the vector of conservative variables after flux integration.
Except for the pressure gradient term of the momentum equation, which is treated during the calculation of numerical fluxes in Athena++, all source terms in the continuity, momentum, and energy equations can be added explicitly with an operator-splitting method, following the standard Athena++ approach. However, examining the source term vector in eq. (30), all of the components in the source term $\mathbf{S}$ can be integrated analytically. Thus, instead of following the standard explicit approach, we incorporate the source terms by updating the conservative variables with the analytical solutions at $t^{n+1}$, namely
@@
with $J^{n}\equiv\mathcal{R}^{n}\mathcal{R}^{n}L_{z}^{n}$ is the Jacobian determinant in the local coordinate system, and
@@
is the kinetic energy density of relative motion at $t^{n+1}$. This approach of analytical integration of the source term significantly enhances the numerical accuracy, and ensures exactly the conservation laws (see §4.3). The method we use for integrating the source term is first-order in time. While higher-order methods have been explored, they are not as efficient within the current framework. We further discuss the impact of the treatment of the source term on numerical accuracy in §4.4.
The flux integration (or the time-averaged numerical fluxes) mentioned above are computed depending on different time integrating methods (see §3.2.3 of Stone et al., 2020, for details). To explain the core of the algorithm, we present hereafter flux calculation in the $x$ direction only as an example, dropping the subscripts $j$ and $k$ for simplicity. For a time integrator with a stage number of $N_{stage}$ in each time step, for each stage $s=0,...N_{stage-1}$, $\mathbf{q}_{i}^{n}$ is updated with fluxes $\mathbf{f}_{i-1/2}^{s}$ and $\mathbf{f}_{i-1/2}^{s}$, which are computed with a Riemann solver. In Athena++, various solvers are implemented, including Roe’s linearized solver (Roe, 1981), as well as the HLL approximated Riemann solvers (Harten et al., 1983). These approximate Riemann solvers provide an estimate of the fluxes. As discussed above, pressure gradient terms in the momentum equation are integrated during the calculation of fluxes in Athena++. Since the corresponding terms in the local model (the right-hand side of Eqs. 68) can be interpreted as an anisotropic sound speed, we modify the Riemann solver, which incorporates the anisotropic effective sound speed by rescaling the wave speed during flux calculation. We apply this modification in Roe’s solver, HLLE solver and HLLC solver, the latter being recommended for adiabatic hydrodynamics. The basic introduction of these Riemann solvers can be found in Toro (1999), so will not be introduced in detail.
For the flux $\mathbf{f}_{i-1/2}^{s}$ at the cell interface $x_{i-1/2}$, the Roe’s solver and HLLE solver provide estimated flux $\mathcal{F}_{i-1/2}^{\mathrm{Roe}}$ and $\mathcal{F}_{i-1/2}^{\mathrm{HLLE}}$ separately. The Roe’s flux is
@@
The HLLE flux at the interface $x_{i-1/2}$ is
@@
In both cases, the subscripts $L$ and $R$ denote the left- and right-states of each variable at the interface, and $\mathbf{f}_{L,i-1/2}=\mathbf{f}\left(\mathbf{q}_{L,i-1/2}\right)$ and $\mathbf{f}_{R,i-1/2}=\mathbf{f}\left(\mathbf{q}_{R,i-1/2}\right)$ are the fluxes evaluated using the left- and right-states of the conserved variables.
@@
where the $\mathscr{L}^{\alpha}$ and $\mathscr{R}^{\alpha}$ are the $\alpha$-th left and right eigenvectors of Roe’s matrix that corresponds to $\lambda^{\alpha}$. $\mathscr{L}^{\alpha}$ and $\mathscr{R}^{\alpha}$ are also rows and columns of the left- and right-eigenmatrices $\mathscr{L}$ and $\mathscr{R}$. For isothermal hydrodynamics, the eigenvalues of Roe’s matrix are
@@
where $C$ is the (effective) sound speed. The left- and right-eigenmatrice
$\mathscr{L}=\left[\begin{array}[]{cccc}\left(1+v_{x}/C\right)/2&-1/(2C)&0&0\\ -v_{y}&0&1&0\\ -v_{z}&0&0&1\\ \left(1-v_{x}/C\right)/2&1/(2C)&0&0\end{array}\right],$ (53)
@@
Figure 2: Test case of horizontal shear flows. Upper panel: 3D visualization of the initial spatial distribution of the shear flow in our simulation. Arrows mark the direction of shear flow, and the total flow velocity magnitude is mapped to colors (see the colorbar). Middle panels: red solid line shows $v_{x}/v_{x0}$ vs. R, i.e. normalized shear velocity as a function of global radius of the local box, showing growth of shear velocity as collapse proceeds, closely matches the analytical expectation (black dashed lines). The result of $v_{y}$ is identical to that of $v_{x}$ and is not shown here. Lower panel: time evolution of relative error $\mathrm{err}(v_{x})$.
Here, we rescale the effective sound speed with a scaling factor $l$, so that $C=c_{\mathrm{s}}/l$, with $c_{\mathrm{s}}$ the isothermal or adiabatic sound speed. On the $x$ and $y$ directions, we choose $l=\mathcal{R}$ so that $C=c_{\mathrm{s},\mathrm{H}}$; on the $z$ direction, $l=L_{z}$ so we have $C=c_{\mathrm{s},\mathrm{V}}$. Note that when $l=1$, we return to the original eigenvalues and eigenvectors as given in Stone et al. (2008). For adiabatic hydrodynamics, the eigenvalues and the eigenmatrice are a bit more complicated than the isothermal case as the (modified) energy equation is also solved. The corresponding quantities are shown in App. A .

### v2

@@
We implement a local model for a spherical collapsing or expanding gas cloud in the Athena++ magnetohydrodynamic code. This local model consists of a Cartesian periodic box with time-dependent geometry. We present a series of benchmark test problems, including nonlinear solutions and linear perturbations of the local model, confirming the code’s desired performance. During a spherical collapse, a horizontal shear flow is amplified, corresponding to angular momentum conservation of zonal flows in the global problem; wave speed and the amplitude of sound waves increase in the local frame, due to the reduction in the characteristic length scale of the box, which can lead to an anisotropic effective sound speed in the local box. Our code conserves both mass and momentum-to-machine precision. This numerical implementation of the local model has potential applications to the study of local physics and hydrodynamic instabilities during protostellar collapse, providing a powerful framework for better understanding the earliest stages of star and planet formation.
@@ 1 Introduction
Star formation has been studied over decades both theoretically and numerically on large spatial scales, the key motivation being to determine how global redistribution of angular momentum results in the formation of stars, disks, or jets and set the initial conditions for planet formation (see Larson 1969; Penston 1969 and e.g., Pattle et al. 2023; Pineda et al. 2023; Tsukamoto et al. 2023 for recent reviews). The field has been stimulated by instruments capable of resolving the largest scales of these young stellar systems in ever-finer detail in several wavelength domains, such as the Atacama Large Millimeter Array (ALMA), the Very Large Telescope (VLT) or the James Webb Space Telescope (JWST). The importance of these questions has perhaps overshadowed that of the physical processes taking place simultaneously on a small scale. One question, for example, is whether it is possible for dust to start concentrating as soon as the cloud collapses, thereby promoting highly efficient planet formation later on.
@@
The local box formalism provides a simple treatment of the hydrodynamics of a spherically symmetric system undergoing contraction. The global effects of curvature and acceleration due to collapse are consistently incorporated into the local treatment of the problem by source terms that can be time-dependent, while allowing for the adoption of periodic boundaries as the system is spatially homogeneous. Numerical simulations using the local approximation approach play an important role in the study of local gas dynamics and instabilities in astrophysics. The best-known example is the rediscovery of the magnetorotational instability (MRI) with the local shear box approximation, which was first applied to magnetohydrodynamic (MHD) simulations by Hawley et al. (1995). Local models have been used for (isotropic) cosmological collapses (Robertson & Goldreich, 2012). Anisotropy in collapse and expansion was taken into account in the study of starless molecular core collapse (Toci et al., 2018), solar winds (Velli et al., 1992; Grappin et al., 1993; Grappin & Velli, 1996; Tenerani & Velli, 2017; Shi et al., 2020; Squire et al., 2020) (among which Tenerani & Velli (2017) also consider nonuniform expansion), anisotropy-driven plasma kinetic instabilities (Sironi & Narayan, 2015; Bott et al., 2021), and the cosmic ray pressure anisotropy instability (Sun & Bai, 2023).
@@
Recently, Lynch & Laibe (2023) (hereafter LL23) developed a new local model for the spherical collapse or expansion problem, allowing for a nonuniform collapse or expansion. This model considers a radially varying flow, as in Tenerani & Velli (2017), but also allows for the time-dependence of the background flow. It also uses a different treatment of pressure gradient than the expanding box models, in an approach that is more similar to the distorted shearing box models of Ogilvie & Latter (2013); Ogilvie & Barker (2014). This avoids spurious instabilities from the inclusion of formally subdominant terms (Latter & Papaloizou, 2017), allowing for the application of periodic or shear-periodic boundary conditions with a Cartesian-like geometry. In this study, we incorporate LL23’s model into Athena++, a high-order Godunov MHD code (Stone et al., 2020). Other local models have already been implemented in Athena++, including Squire et al. (2020) for the (uniformly) expanding box, and MHD-PIC (Sun & Bai, 2023), also allowing for an anisotropic collapse or expansion. However, most of these previous numerical implementations focus on the study of MHD properties and magnetized waves, in the context of stellar winds or cosmic rays, and the hydrodynamic properties of the expanding or collapsing box have not been comprehensively tested so far. Here, we aim to incorporate the local model of LL23, with benchmark tests to thoroughly investigate and validate the hydrodynamic properties therein.
@@
This article is organized as follows. We first describe the formulation of the local model for the collapsing box in Sect. 2. In Sect. 3, we describe the numerical implementation, especially the modification of the Riemann solver in order to incorporate the local model. We then conduct a series of test cases including nonlinear solutions and linear perturbations in Sect. 4, as well as numerical studies. We further discuss possible applications and perspectives, and give future perspectives in Sect. 5. We summarize and conclude in Sect. 6.
@@
Figure 1: Cartoon illustration of the local collapsing box model. On the left, the global geometry depicts a domain moving radially due to background flow. The volume of the domain varies as it approaches the center of the spherical cloud, leading to changes in density. A compressing domain leads to a density increase, as is shown by the progressively darker blue color in the sequence. The horizontal domain size is proportional to $\mathcal{R}\left(t\right)$, establishing the horizontal length scale in the local model (middle panel). The vertical length scale, $L_{z}\left(t\right)$, is determined by the distance between the domain’s inner and outer radial boundaries, which is set by the background flow. Inside each box, a waveform signal illustrates how certain physical quantities evolve in the local model. The width of the waveform represents quantities that remains fixed in the local frame (e.g., wavelength). The amplitude of the waveform represents specific physical quantities that are expected to be constant in the global geometry throughout the collapse (e.g., angular momentum, isothermal physical sound speed). In the local reference frame (right), the corresponding quantities such as the shear flow velocity or the effective sound speed are amplified as the collapse proceeds and the length scale shrinks in the domain, while quantities such as the wavelength remain the same. This phenomenon is visually demonstrated by the vertically stretched waveforms in the sequence. For a more accurate and detailed understanding of this process, one should refer to the equations in §2 .
@@ 2 Equations of the local model
We started by considering an axisymmetric spherical gas cloud collapsing or expanding with a purely radial velocity, $U_{bg}(r,t)$, that varies with both radius, $r$, and time, $t$. We then considered a local neighborhood of a reference point located at a radius, $\mathcal{R}(t)$, in the spherical gas cloud, described by a local Cartesian coordinate system comoving with the background flow. The local coordinates are given by
@@
where $\phi$ and $\theta$ are the longitude and latitude on the reference sphere of radius, $\mathcal{R}$, and $r-\mathcal{R}(t)$ is the local radial distance relative to the global, $\mathcal{R}(t)$, of the reference frame. We note that $x$, $y$, $z$ have dimensions of angles in this formalism, and that they denote two azimuthal and radial directions, respectively. $L_{z}$ is a characteristic vertical length scale that is set by the background flow profile, which can be approximated by
@@
The flow within the local box model represents local, nonlinear perturbations to the background flow. Following the derivation in Sect. 2 of LL23, in the local frame, the continuity, momentum, and energy equations are given in the stretched Cartesian coordinate system by
@@
where $\rho$ is the density of the gas, $v_{i}$ is the relative velocity field in the local frame, [^1] The velocities here are the contravariant components of the velocity. In this article, we use subscript indices to denote the spatial dimensions (e.g., $v_{i}$) for convenience, as opposed to superscript indices (e.g., $v^{i}$) as is conventionally used. We also apply Einstein summation throughout this paper, where repeated indices (subscript or superscript) are implicitly summed over. $p$ is the pressure of the gas, and $\mathcal{E}_{rell}$ is the energy density of the relative motion in this system, (re-)defined as
$\mathcal{E}_{\text{rell }}=\frac{1}{2}L_{z}^{2}v_{z}^{2}+\frac{1}{2}\mathcal{R}^{2}(t)\left(v_{x}^{2}+v_{y}^{2}\right)+\varepsilon,$ (10)
which is the sum of the kinetic energy density of relative motion in the local frame, $\mathcal{E}_{k,\text{rell }}\equiv\frac{1}{2}L_{z}^{2}v_{z}^{2}+\frac{1}{2}\mathcal{R}^{2}(t)\left(v_{x}^{2}+v_{y}^{2}\right)$, and the specific internal energy, $\varepsilon$, determined by the equation of state (EoS). For an isothermal EoS, $p=\rho c_{s}^{2}$, where $c_{s}$ is the sound speed, and the energy equation (9) is dropped from the system. For an adiabatic EoS, $\varepsilon=\frac{p}{\rho(\gamma-1)}$ , where $\gamma$ is the ratio of specific heats, assumed to be constant.
The right-hand side of Eqs. 68 can be interpreted as an anisotropic sound speed in the local coordinate. We define two rescaled sound speeds in the horizontal and vertical directions separately: $c_{\mathrm{s},\mathrm{H}}\equiv c_{\mathrm{s}}/\mathcal{R}$ and $c_{\mathrm{s},\mathrm{V}}\equiv c_{\mathrm{s}}/L_{z}$.
Eqs. 59 can be written in a compact form,
@@
where the vector, $\mathbf{q}$, of conservative variables is defined according to
@@
$\mathbf{f},\mathbf{g}$, and $\mathbf{h}$ are the fluxes of the associated conservative variables in the $x$, $y$, and $z$ directions, respectively:
@@
The vector of source terms, denoted as $\mathbf{S}$, is
@@
Again, for an isothermal EoS, the fifth component of the vectors $\mathbf{q}$, $\mathbf{f}$, $\mathbf{g}$, $\mathbf{h}$, and $\mathbf{S}$ are dropped, as the energy equation is not solved for this case.
A more intuitive illustration of the local model geometry is presented in Fig. 1. As the collapse proceeds, the local domain is compressed in volume, leading to an increase in gas density. This is reflected in the source term in the continuity equation (5). The angular momentum conservation law leads to amplification of horizontal flow velocity in the local frame during the collapse, which is represented by the source terms in the momentum equations on the horizontal directions (Eqs. 6 and 7). These features along with other characteristic behaviors of the local model are further illustrated and tested in Sect. 4.
@@ 3 Numerical implementation
We next translated the local model of § 2 into a Godunov scheme, and implemented it in the numerical code Athena++. Finite volume discretization of the compact form of the hydrodynamic equations Eq.(11) gives the time integration step as a set of ordinary differential equations,
@@
where subscript indices $(i,j,k)$ denote numerical cells centered at positions $(x_{i},y_{j},z_{k})$, superscripts, $n$, denotes the index for discrete time, $\delta x$, $\delta y$, and $\delta z$ are discrete grid sizes, and $\delta t$ is the time step. The vectors $\mathbf{q}_{i,j,k}^{n}$ and $\mathbf{S}_{i,j,k}^{n}$ correspond to volume-averaged $\mathbf{q}$ and $\mathbf{S}$ over the numerical cells. The vectors $\mathbf{f}_{i+1/2,j,k}^{n+1/2}$, $\mathbf{g}_{i+1/2,j,k}^{n+1/2}$, and $\mathbf{h}_{i+1/2,j,k}^{n+1/2}$ correspond to time-averaged $\mathbf{f}$, $\mathbf{g}$, and $\mathbf{h}$, respectively. The half-integer subscripts on the flux vectors refer to the edges of computational cells, while the half-integer superscripts denote the time average between $t^{n}$ and $t^{n+1}$. $\mathbf{\tilde{q}}_{i,j,k}^{n}$ denotes the vector of conservative variables after flux integration.
Except for the pressure gradient term of the momentum equation, which was treated during the calculation of numerical fluxes in Athena++, all source terms in the continuity, momentum, and energy equations can be added explicitly with an operator-splitting method, following the standard Athena++ approach. However, examining the source term vector in eq. (30), all of the components in the source term, $\mathbf{S}$, can be integrated analytically. Thus, instead of following the standard explicit approach, we incorporated the source terms by updating the conservative variables with the analytical solutions at $t^{n+1}$; namely,
@@
with $J^{n}\equiv\mathcal{R}^{n}\mathcal{R}^{n}L_{z}^{n}$ the Jacobian determinant in the local coordinate system, and
@@
the kinetic energy density of relative motion at $t^{n+1}$. This approach of analytical integration of the source term significantly enhances the numerical accuracy, and ensures that the conservation laws are fully obeyed (see Sect. 4.3). The method we used to integrate the source term is first-order in time. While higher-order methods have been explored, they are not as efficient within the current framework. We further discuss the impact of the treatment of the source term on numerical accuracy in Sect. 4.4.
The flux integration (or the time-averaged numerical fluxes) mentioned above were computed depending on different time-integrating methods (see §3.2.3 of Stone et al., 2020, for details). To explain the core of the algorithm, we present hereafter the flux calculation in the $x$ direction only as an example, dropping the subscripts $j$ and $k$ for simplicity. For a time integrator with a stage number of $N_{stage}$ in each time step, for each stage $s=0,...N_{stage-1}$, $\mathbf{q}_{i}^{n}$ was updated with fluxes $\mathbf{f}_{i-1/2}^{s}$ and $\mathbf{f}_{i-1/2}^{s}$, which were computed with a Riemann solver. In Athena++, various solvers were implemented, including Roe’s linearized solver (Roe, 1981), as well as the Harten-Lax-van Leer (HLL) approximated Riemann solvers (Harten et al., 1983). These approximate Riemann solvers provide an estimate of the fluxes. As was discussed above, pressure gradient terms in the momentum equation were integrated during the calculation of fluxes in Athena++. Since the corresponding terms in the local model (the right-hand side of Eqs. 68) can be interpreted as an anisotropic sound speed, we modified the Riemann solver, which incorporates the anisotropic effective sound speed by rescaling the wave speed during flux calculation. We applied this modification in Roe’s solver, the Harten-Lax-van Leer-Einfeldt (HLLE) solver, and the Harten-Lax-van Leer-Contact (HLLC) solver, the latter being recommended for adiabatic hydrodynamics. The basic introduction of these Riemann solvers can be found in Toro (1999), so will not be introduced in detail.
For the flux, $\mathbf{f}_{i-1/2}^{s}$, at the cell interface, $x_{i-1/2}$, the Roe’s solver and HLLE solver provide the estimated flux, $\mathcal{F}_{i-1/2}^{\mathrm{Roe}}$ and $\mathcal{F}_{i-1/2}^{\mathrm{HLLE}}$, separately. The Roe’s flux is
@@
The HLLE flux at the interface, $x_{i-1/2}$, is
@@
In both cases, the subscripts $L$ and $R$ denote the left and right states of each variable at the interface, and $\mathbf{f}_{L,i-1/2}=\mathbf{f}\left(\mathbf{q}_{L,i-1/2}\right)$ and $\mathbf{f}_{R,i-1/2}=\mathbf{f}\left(\mathbf{q}_{R,i-1/2}\right)$ are the fluxes evaluated using the left and right states of the conserved variables.
@@
where the $\mathscr{L}^{\alpha}$ and $\mathscr{R}^{\alpha}$ are the $\alpha$-th left and right eigenvectors of Roe’s matrix, which corresponds to $\lambda^{\alpha}$. $\mathscr{L}^{\alpha}$ and $\mathscr{R}^{\alpha}$ are also rows and columns of the left and right eigenmatrices, $\mathscr{L}$ and $\mathscr{R}$. For isothermal hydrodynamics, the eigenvalues of Roe’s matrix are
@@
where $C$ is the (effective) sound speed. The left and right eigenmatrices are
$\mathscr{L}=\left[\begin{array}[]{cccc}\left(1+v_{x}/C\right)/2&-1/(2C)&0&0\\ -v_{y}&0&1&0\\ -v_{z}&0&0&1\\ \left(1-v_{x}/C\right)/2&1/(2C)&0&0\end{array}\right]$ (53)
and
@@
Figure 2: Test case of horizontal shear flows. Upper panel: 3D visualization of the initial spatial distribution of the shear flow in our simulation. Arrows mark the direction of shear flow, and the total flow velocity magnitude is mapped to colors (see the color bar). Middle panels: Solid red line shows $v_{x}/v_{x0}$ vs. R, i.e., normalized shear velocity as a function of global radius of the local box, showing growth of shear velocity as collapse proceeds, closely matches the analytical expectation (dashed black lines). The result of $v_{y}$ is identical to that of $v_{x}$ and is not shown here. Lower panel: Time evolution of relative error $\mathrm{err}(v_{x})$.
Here, we rescaled the effective sound speed with a scaling factor, $l$, so that $C=c_{\mathrm{s}}/l$, with $c_{\mathrm{s}}$ the isothermal or adiabatic sound speed. In the $x$ and $y$ directions, we chose $l=\mathcal{R}$ so that $C=c_{\mathrm{s},\mathrm{H}}$; in the $z$ direction, $l=L_{z}$ so we have $C=c_{\mathrm{s},\mathrm{V}}$. We note that when $l=1$, we returned to the original eigenvalues and eigenvectors given in Stone et al. (2008). For adiabatic hydrodynamics, the eigenvalues and the eigenmatrice are a bit more complicated than the isothermal case as the (modified) energy equation is also solved. The corresponding quantities are shown in App. A .